OpenStudy (anonymous):

What are the next 2 terms in the following sequences? -1,-3,-5,-7,...

OpenStudy (anonymous):

u see the pattern is decreasing by -2 each time? -1-2=-3 -3-2=-5 so for the next two terms we do this. -7-2=-9 -9-2=-11